# Landing with the Score: Riemannian Optimization through Denoising

### v1

@@ 2 Preliminaries
Here we introduce some preliminary notation and concepts to state our results. We refer the reader to the Appendices A and C for more details.
@@ 2.1 Manifolds and distance functions
Let $\mathcal{M}\subseteq\mathbb{R}^{d}$ be a $k$-dimensional embedded compact $C^{2}$-submanifold without boundary. For any point $p\in\mathcal{M}$ we denote by $\operatorname{T}_{p}\mathcal{M}\subseteq\mathbb{R}^{d}$ and $\operatorname{N}_{p}\mathcal{M}\subseteq\mathbb{R}^{d}$ the tangent and normal spaces of $\mathcal{M}$ at $p$, respectively, and their orthogonal projections by $\operatorname{P}_{\operatorname{T}_{p}\mathcal{M}}$ and $\operatorname{P}_{\operatorname{N}_{p}\mathcal{M}}$. The squared distance function is defined by $\operatorname{d}(x)=\inf_{p\in\mathcal{M}}\frac{1}{2}\lVert x-p\rVert^{2}$. For a $C^{2}$-submanifold, then there exists a radius $\tau_{\mathcal{M}}>0$ such that every point in $x\in\mathcal{T}(\tau)=\bigcup_{p\in\mathcal{M}}B_{\tau}(p)$ has a unique projection $\pi(x)\in\mathcal{M}$ such that $\frac{1}{2}\lVert x-\pi(x)\rVert^{2}=\operatorname{d}(x)$ and $x-\pi(x)\in\operatorname{N}_{\pi(x)}\mathcal{M}$. Sets of the form $\mathcal{T}(\tau)$ will be called tubular neighborhoods of $\mathcal{M}$ with closure denoted by $\overline{\mathcal{T}}(\tau)$. One can prove that $\operatorname{d}$ and $\pi$ are differentiable on $\mathcal{T}(\tau_{\mathcal{M}})$ with $\operatorname{d}^{\prime}(x)=x-\pi(x)$ and $\pi^{\prime}(x)$ given explicitly in Appendix A. The maximal principal curvature of a manifold will be denoted by $\kappa_{\mathcal{M}}$. It always holds that $\tau_{\mathcal{M}}\leq 1/\kappa_{\mathcal{M}}$. Finally, for a function $f:\mathcal{M}\to\mathbb{R}$ we denote the Riemannian gradient by $\operatorname{grad}_{\mathcal{M}}f(p)$, which in the case of $f:\mathbb{R}^{d}\to\mathbb{R}$ is given by $\operatorname{grad}_{\mathcal{M}}f(p)=\operatorname{P}_{\operatorname{T}_{p}\mathcal{M}}\nabla f(p)$. We write
@@ 2.2 The Stein score function and score-based diffusion models
For a Borel probability measure we denote its Gaussian blurring by $p_{\sigma}=\mathcal{N}(0,\sigma^{2}I)\ast\mu$ for $\sigma>0$ with $p_{0}=\mu$. The score function $\nabla\log p_{\sigma}$ of $p_{\sigma}$ and its Jacobian allow for the following interpretation (see Jaffer & Gupta (1972), also Appendix B):
@@
where $\ell_{\sigma}$ is the link function (3) and $\nu_{x,\sigma}$ is the posterior distribution observing $x$ under the noise model $p_{\sigma}$ and prior $\mu$. The representation for $\mathbb{E}\nu_{x,\sigma}$ has also been known under the name of Tweedie’s formula Robbins (1992); Efron (2011). The score function has gained recent attention due to its use in score-based diffusion models in the field of generative modelling. Specifically in the so-called variance exploding (VE) diffusion scheme one seeks to learn $\nabla\log p_{\sigma}$ for different noise scales $\sigma$ via a neural network $\texttt{s}(\cdot,\sigma)$ by minimizing the conditional score matching loss $L_{\operatorname{CSM}}(\texttt{s}(\cdot,\sigma))$, which attains its unique minimum in $\texttt{s}(\cdot,\sigma)=\nabla\log p_{\sigma}$. Then $\texttt{s}(\cdot,\sigma)$ is used for sampling from $\mu$ by following a particular reverse-time SDE or ODE flow in the noise scale $\sigma$ (see Appendix C for more details).
@@ 3 Score as retraction and tangent space projection
The proof is deferred to Appendix D and is based on a careful non-asymptotic estimate of the Laplace integral method. As a consequence and together with fact that $\pi^{\prime}(x)$ coincides with $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ for $x\in\mathcal{M}$ (see Appendix A) we obtain the following result.
@@
Thus, in theory, a well-trained diffusion model score $\texttt{s}(\cdot,\sigma)$ and its Jacobian $\texttt{s}^{\prime}(\cdot,\sigma)$ allow us to approximate the closest-point projection $\pi(x)$ and the tangent space projection $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ as $\sigma\to 0$ arbitrarily well via its Tweedie score $x+\sigma^{2}\texttt{s}(x,\sigma)$ and its Jacobian, respectively.
@@ 4 Denoising Riemannian gradient flow with landing
In this section we show how to use the score function from Section 3 for Riemannian optimization of (1) for some smooth $f\in C^{1}(\mathbb{R}^{d})$. Assuming that we have access to a sufficiently accurate estimate of the Tweedie score in form of a vector function $s\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$ such that $s(x)=x+\sigma^{2}\texttt{s}(x,\sigma)$ and
$\lVert s(x)-\pi_{\sigma}(x)\rVert\leq\epsilon\text{\ and\ }\lVert s^{\prime}(x)-P_{\sigma}(x)\rVert\leq\epsilon\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$ (7)
@@
$\dot{x}$ $=-s^{\prime}(x)\nabla f(s(x))+\eta(s(x)-x)\,.$ (8)
In the exact case $s(x)=\pi_{\sigma}(x)$ and $s^{\prime}(x)=P_{\sigma}(x)$ (i.e. $\epsilon=0$ in (7)), flow (8) is the gradient flow
@@
and the dynamics in (9) consists of two parts: An approximate projection $P_{\sigma}(x)\nabla f(\pi_{\sigma}(x))$ of the gradient $\nabla f(\pi_{\sigma}(x))$ and an approximate landing term $\eta(\pi_{\sigma}(x)-x)$ corresponding to the penalty function $\eta\operatorname{d}_{\sigma}(x)$. In the further case of $\sigma=0$ and $x(0)\in\mathcal{M}$ the flow (9) reduces to the ordinary Riemannian gradient flow, which has been extensively studied Helmke & Moore (2012); Ambrosio et al. (2005). Interestingly, when $\sigma=0$, but only $x(0)\in\mathcal{T}(\tau)$, then (9) reduces to
@@
with $H_{x}^{-1}$ a linear operator on $\operatorname{T}_{\pi(x)}\mathcal{M}$ given in Appendix A. In particular the two terms in (10) belong to $\operatorname{T}_{\pi(x)}\mathcal{M}$ and $\operatorname{N}_{\pi(x)}\mathcal{M}$, respectively, and are orthogonal to each other, implying that the distance between $x$ and $\mathcal{M}$ is non-increasing, which allows for perfect landing on the manifold via similar arguments as in Ablin & Peyré (2022); Schechtman et al. (2023), see Theorem 20 in Appendix E.3. For $\sigma>0$ or $\epsilon>0$, the two summands in (9) and (8) in general not perpendicular to each other and the landing is not exact. However, using Theorem 1 we show the following.
@@
Consider the flow (8) and $\tau\in(0,\tau_{\mathcal{M}})$. Set $C=\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f)$. Suppose that for some $\epsilon>0$ and $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ with $\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\min\{\tau,\frac{2\tau}{1+C/\eta}\}$ the function $s$ satisfies (7). Then for any $x(0)\in\mathcal{T}(\tau)$ the solution $x(t)$ to (8) exists for all $t\geq 0$ and is contained in $\mathcal{T}(\tau)$. Moreover, every accumulation point $x_{*}$ of this flow satisfies
@@
We can evaluate the right hand side of the flow (8) in a single forward-backward pass of the network $s$. Namely, given an input $x$, we compute and store $p=s(x)$ by a forward pass of $s$, while keeping the computational graph of $s(x)$. Then we evaluate $v=\nabla f(p)$ and build the computational graph of $y=\left\langle s(x),v\right\rangle$, while detaching $v$. Finally we backpropagate on $x$ in $y$ to obtain $s^{\prime}(x)v=s^{\prime}(x)\nabla f(s(x))$.
@@ 5 Denoising Riemannian gradient descent
$x_{k+1}=s(x_{k}-\gamma_{k}s^{\prime}(x_{k})\nabla f(x_{k}))\,$, (12)
which we term the denoising Riemannian gradient descent (DRGD). Here $s$ acts as an approximate retraction and $s$ as an approximate projection onto the tangent space. We obtain the following convergence result for this algorithm.
@@
Suppose that for some $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ and $\epsilon>0$ with $\epsilon^{\prime}:=\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\tau/2$ the function $s$ satisfies (7) and that the step-size $\gamma_{k}$ is constrained by $\gamma_{k}\in[\gamma_{\min},\gamma_{\max}]$ with
@@
Note that both Theorem 3 and Theorem 5 require the rather strong $L^{\infty}$-approximation assumption (7) on the Tweedie score $s$ and its Jacobian. The analysis under a weaker $L^{2}$-bound is out of score for this paper and left for future work.
@@ 6.1 Optimization on orthogonal group $O(n)$
In this section we evaluate flow (8) on a synthetic example. In order to compare the error of our method to classical Riemannian optimization techniques, we consider distributions supported on manifolds that are the focus of study in Absil et al. (2008); Boumal (2023). Specifically we consider the orthogonal group manifold $\mathcal{M}=\operatorname{O}(n)\subseteq\mathbb{R}^{n\times n}$ with $\mu=\operatorname{Vol}_{\mathcal{M}}$ being the uniform volume measure and Brockett’s cost function Helmke & Moore (2012) defined by
@@
where $A,Q\in\mathbb{S}^{n\times n}$ are given. We consider the cases $n\in\{10,20\}$ and assume that we are given a set $\mathcal{D}_{\operatorname{train}}\subseteq\mathcal{M}$ of $N_{\operatorname{data}}=20000$ data points from $\mu$ and train the score function $s$ with denoising score matching (see Appendix C for diffusion models and Appendix G.1.1 for implementation details). In Figure 3 (left) (in Appendix G.1.2 for space reasons) we compare the evolution of the objective value of our approximation of (8) to the exact landing flow (10) for different noise levels $\sigma>0$ and dimension $n$. We observe that we can obtain objective values with cost lower than the best possible point in the training set and that the accuracy improves as $\sigma\to 0$.
@@ 6.2 Reference tracking via data-driven control
Under certain smoothness assumptions on the dynamics $\bm{f}$ and $\bm{g}$ the set $\mathcal{M}_{\operatorname{IO}}$ is a (non-compact) embedded smooth submanifold of $(\mathbb{R}^{n_{u}})^{N_{h}}\times(\mathbb{R}^{n_{y}})^{N_{h}+1}$. The problem (14) is ubiquitous in receding horizon control applications such as model predictive control (MPC) and used for e.g. autonomous driving Vu et al. (2021), motion planning Cohen et al. (2020), optimizing HVAC system energy efficiency Serale et al. (2018) and inventory control Kostić (2009). In many of these applications the dynamics (13) governing the system are not known explicitly. Instead, in data-driven control Dörfler (2023a; b); Markovsky et al. (2023) one assumes that (13) is given implicitly by a finite number of measured input-output trajectories
@@
where the input $\bm{u}$ is persistently exciting Willems et al. (2005), e.g. given by (white) noise Ljung (1999). In particular the so-called system behavior manifold $\mathcal{M}_{\operatorname{IO}}$ Willems & Polderman (1997) is given by samples from a distribution $\mu$ on its in- and outputs and fits precisely into our framework of data-driven Riemannian optimization (1). We test our proposed denoising Riemannian gradient descent on two classical systems from the control domain: The discretized double pendulum system and the unicycle car model LaValle (2006) (see Appendix G.2.1 detailed information on the systems and the particular choice of $\bm{r}$, $R$ and $Q$ in (15)), each on a horizon of $N_{h}=100$. To apply our proposed methods, we train a diffusion model (see Appendix G.2.2 for implementation details) on the measured trajectories $\mathcal{D}_{\operatorname{train}}$ and solve (14) via the denoising Riemannian gradient descent to obtain a solution $(\bm{u}^{*},\bm{y}^{*})$. As initial values we take the trajectories from the training set that minimize the objective cost $f$, i.e. $(\bm{u}_{0},\bm{y}_{0})=\operatorname*{arg\,min}_{(\bm{u},\bm{y})\in\mathcal{D}_{\operatorname{train}}}f(\bm{u},\bm{y})$. Note that, as seen in Section 5, in general the final iterate will not exactly lie on the input-output manifold, i.e. $(\bm{u}^{*},\bm{y}^{*})\notin\mathcal{M}_{\operatorname{IO}}$. To account for this deviation we back-test our generated input trajectory $\bm{u}^{*}$ by implementing it on the true underlying system (13) to obtain the real output $\bm{y}^{\operatorname{true}}$.
Results and discussion: In Figure 4 (in Appendix G.2.3) we depict the evolution of the objective value w.r.t. iteration count and in Figure 2 we depict the final optimizing trajectories $\bm{y}^{*}$ and $\bm{y}^{\operatorname{true}}$. We can observe that the error $\lVert\bm{y}^{*}-\bm{y}^{\operatorname{true}}\rVert$ is small, which shows that $(\bm{u}^{*},\bm{y}^{*})$ is close to the true system behavior $\mathcal{M}_{\operatorname{IO}}$. Moreover, we can see that the trajectory $\bm{y}^{\operatorname{true}}$ tracks the corresponding reference $\bm{r}$ much better than the train set minimum $\bm{y}_{0}$, which shows a generalization capability of our diffusion model. We from Figure 4 (left) that the current objective can depart from the true objective significantly. This is due to the iterates deviating from $\mathcal{M}_{\operatorname{IO}}$. In this example the algorithm (DRGD) recovers and we have found it to be robust w.r.t. moderate deviations from the manifold. Note that we have set a iteration budget of $N_{\operatorname{iter}}=3000$ and $N_{\operatorname{iter}}=2500$, respectively, while the objective is still decreasing. Accelerating the convergence of DRGD is left for future work.

### v2

@@ 2 Preliminaries
Here we introduce some preliminary notation and concepts to state our results. We refer the reader to the Appendices B and D for more details.
@@ 2.1 Manifolds and distance functions
Let $\mathcal{M}\subseteq\mathbb{R}^{d}$ be a $k$-dimensional embedded compact $C^{2}$-submanifold without boundary. For any point $p\in\mathcal{M}$ we denote by $\operatorname{T}_{p}\mathcal{M}\subseteq\mathbb{R}^{d}$ and $\operatorname{N}_{p}\mathcal{M}\subseteq\mathbb{R}^{d}$ the tangent and normal spaces of $\mathcal{M}$ at $p$, respectively, and their orthogonal projections by $\operatorname{P}_{\operatorname{T}_{p}\mathcal{M}}$ and $\operatorname{P}_{\operatorname{N}_{p}\mathcal{M}}$. For a $C^{2}$-submanifold, then there exists a radius $\tau_{\mathcal{M}}>0$ such that every point in $x\in\mathcal{T}(\tau_{\mathcal{M}})$, where
$\mathcal{T}(\tau):=\{x\in\mathbb{R}^{d}\mid\operatorname{dist}(x,\mathcal{M})<\tau\}\,,\quad\operatorname{dist}(x,\mathcal{M})=\inf_{p\in\mathcal{M}}\lVert x-p\rVert\,$,
has a unique projection $\pi(x)\in\mathcal{M}$. Sets of the form $\mathcal{T}(\tau)$ with $\tau\in(0,\tau_{\mathcal{M}})$ are called tubular neighborhoods of $\mathcal{M}$. The squared distance function is defined by $\operatorname{d}(x)=\frac{1}{2}\operatorname{dist}(x,\mathcal{M})^{2}$ and $\operatorname{d}$ and $\pi$ are both differentiable on $\mathcal{T}(\tau_{\mathcal{M}})$ with $\frac{1}{2}\lVert x-\pi(x)\rVert^{2}=\operatorname{d}(x)$ and $x-\pi(x)\in\operatorname{N}_{\pi(x)}\mathcal{M}$ and $\operatorname{d}^{\prime}(x)=x-\pi(x)$ and $\pi^{\prime}(x)$ given explicitly in Appendix B. Finally, for a function $f:\mathcal{M}\to\mathbb{R}$ we denote the Riemannian gradient by $\operatorname{grad}_{\mathcal{M}}f(p)$, which in the case of $f:\mathbb{R}^{d}\to\mathbb{R}$ is given by $\operatorname{grad}_{\mathcal{M}}f(p)=\operatorname{P}_{\operatorname{T}_{p}\mathcal{M}}\nabla f(p)$.
@@ 2.2 The Stein score function and score-based diffusion models
For a Borel probability measure we denote its Gaussian blurring by $p_{\sigma}=\mathcal{N}(0,\sigma^{2}I)\ast\mu$ for $\sigma>0$ with $p_{0}=\mu$. The score function $\nabla\log p_{\sigma}$ of $p_{\sigma}$ and its Jacobian allow for the following interpretation (see Jaffer and Gupta (1972), also Appendix C):
@@
where $\ell_{\sigma}$ is the link function (3) and $\nu_{x,\sigma}$ is the posterior distribution observing $x$ under the noise model $p_{\sigma}$ and prior $\mu$. The representation for $\mathbb{E}\nu_{x,\sigma}$ has also been known under the name of Tweedie’s formula Robbins (1992); Efron (2011). The score function has gained recent attention due to its use in score-based diffusion models in the field of generative modelling. Specifically in the so-called variance exploding (VE) diffusion scheme one seeks to learn $\nabla\log p_{\sigma}$ for different noise scales $\sigma$ via a neural network $\texttt{s}(\cdot,\sigma)$ by minimizing the conditional score matching loss $L_{\operatorname{CSM}}(\texttt{s}(\cdot,\sigma))$, which attains its unique minimum in $\texttt{s}(\cdot,\sigma)=\nabla\log p_{\sigma}$. Then $\texttt{s}(\cdot,\sigma)$ is used for sampling from $\mu$ by following a particular reverse-time SDE or ODE flow in the noise scale $\sigma$ (see Appendix D for more details).
@@ 3 Score as retraction and tangent space projection
The proof is deferred to Appendix E and is based on a careful non-asymptotic estimate of the Laplace integral method. As a consequence and together with fact that $\pi^{\prime}(x)$ coincides with $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ for $x\in\mathcal{M}$ (see Appendix B) we obtain the following result.
@@
Thus, in theory, a well-trained diffusion model score $\texttt{s}(\cdot,\sigma)$ and its Jacobian $\texttt{s}^{\prime}(\cdot,\sigma)$ allow us to approximate the closest-point projection $\pi(x)$ and the tangent space projection $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ as $\sigma\to 0$ arbitrarily well via the operator $\texttt{v}(x)=x+\sigma^{2}\texttt{s}(x,\sigma)$ and its Jacobian, respectively.
@@ 4 Denoising Riemannian gradient flow with landing
In this section we show how to use the score function from Section 3 for Riemannian optimization of (1) for some smooth $f\in C^{1}(\mathbb{R}^{d})$. Assuming that we have access to a sufficiently accurate estimate of the score in form of a vector function $\texttt{v}\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$ such that $\texttt{v}(x)=x+\sigma^{2}\texttt{s}(x,\sigma)$ and
$\lVert\texttt{v}(x)-\pi_{\sigma}(x)\rVert\leq\epsilon\text{\ and\ }\lVert\texttt{v}^{\prime}(x)-P_{\sigma}(x)\rVert\leq\epsilon\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$ (7)
@@
$\dot{x}$ $=-\texttt{v}^{\prime}(x)\nabla f(\texttt{v}(x))+\eta(\texttt{v}(x)-x)\,.$ (8)
In the exact case $\texttt{v}(x)=\pi_{\sigma}(x)$ and $\texttt{v}^{\prime}(x)=P_{\sigma}(x)$ (i.e. $\epsilon=0$ in (7)), flow (8) is the gradient flow
@@
and the dynamics in (9) consists of two parts: An approximate projection $P_{\sigma}(x)\nabla f(\pi_{\sigma}(x))$ of the gradient $\nabla f(\pi_{\sigma}(x))$ and an approximate landing term $\eta(\pi_{\sigma}(x)-x)$ corresponding to the penalty function $\eta\operatorname{d}_{\sigma}(x)$. In the further case of $\sigma=0$ and $x(0)\in\mathcal{M}$ the flow (9) reduces to the ordinary Riemannian gradient flow, which has been extensively studied Helmke and Moore (2012); Ambrosio et al. (2005). Interestingly, when $\sigma=0$, but only $x(0)\in\mathcal{T}(\tau)$, then (9) reduces to
@@
with $H_{x}^{-1}$ a linear operator on $\operatorname{T}_{\pi(x)}\mathcal{M}$ given in Appendix B. In particular the two terms in (10) belong to $\operatorname{T}_{\pi(x)}\mathcal{M}$ and $\operatorname{N}_{\pi(x)}\mathcal{M}$, respectively, and are orthogonal to each other, implying that the distance between $x$ and $\mathcal{M}$ is non-increasing, which allows for perfect landing on the manifold via similar arguments as in Ablin and Peyré (2022); Schechtman et al. (2023), see Theorem 22 in Appendix F.3. For $\sigma>0$ or $\epsilon>0$, the two summands in (9) and (8) in general not perpendicular to each other and the landing is not exact. However, using Theorem 1 we show the following.
@@
Consider the flow (8) and $\tau\in(0,\tau_{\mathcal{M}})$. Set $C=\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f)$. Suppose that for some $\epsilon>0$ and $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ with $\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\min\{\tau,\frac{2\tau}{1+C/\eta}\}$ the function v satisfies (7). Then for any $x(0)\in\mathcal{T}(\tau)$ the solution $x(t)$ to (8) exists for all $t\geq 0$ and is contained in $\mathcal{T}(\tau)$. Moreover, every accumulation point $x_{*}$ of this flow satisfies
@@
We can evaluate the right hand side of the flow (8) in a single forward-backward pass of the network v. Namely, given an input $x$, we compute and store $p=\texttt{v}(x)$ by a forward pass of v, while keeping the computational graph of $\texttt{v}(x)$. Then we evaluate $g=\nabla f(p)$ and build the computational graph of $y=\left\langle\texttt{v}(x),g\right\rangle$, while detaching $g$. Finally we backpropagate on $x$ in $y$ to obtain $\texttt{v}^{\prime}(x)g=\texttt{v}^{\prime}(x)\nabla f(\texttt{v}(x))$.
@@ 5 Denoising Riemannian gradient descent
$x_{k+1}=\texttt{v}(x_{k}-\gamma_{k}\texttt{v}^{\prime}(x_{k})\nabla f(x_{k}))\,$, (12)
which we term the denoising Riemannian gradient descent (DRGD). Here v acts as an approximate retraction and $\texttt{v}^{\prime}$ as an approximate projection onto the tangent space. We obtain the following convergence result for this algorithm.
@@
Suppose that for some $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ and $\epsilon>0$ with $\epsilon^{\prime}:=\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\tau/2$ the function v satisfies (7) and that the step-size $\gamma_{k}$ is constrained by $\gamma_{k}\in[\gamma_{\min},\gamma_{\max}]$ with
@@
If we set $\epsilon=0$ and $\sigma=0$, we recover, up to constants, the classical result on iterates of the Riemannian gradient descent with known manifold $\mathcal{M}$ and non-convex objective $f$ with Lipschitz gradient (Boumal, 2023, Corollary 4.9) (see Appendix A.3).
Remark 7.
Note that both Theorem 3 and Theorem 5 require the rather strong $L^{\infty}$-approximation assumption (7) on v and its Jacobian. The analysis under a weaker $L^{2}$-bound is out of score for this paper and left for future work.
@@ 6.1 Optimization on orthogonal group $O(n)$
In this section we evaluate flow (8) on a synthetic example. In order to compare the error of our method to classical Riemannian optimization techniques, we consider distributions supported on manifolds that are the focus of study in Absil et al. (2008); Boumal (2023). Specifically we consider the orthogonal group manifold $\mathcal{M}=\operatorname{O}(n)\subseteq\mathbb{R}^{n\times n}$ with $\mu=\operatorname{Vol}_{\mathcal{M}}$ being the uniform volume measure and Brockett’s cost function Helmke and Moore (2012) defined by
@@
where $A,Q\in\mathbb{S}^{n\times n}$ are given. We consider the cases $n\in\{10,20\}$ and assume that we are given a set $\mathcal{D}_{\operatorname{train}}\subseteq\mathcal{M}$ of $N_{\operatorname{data}}=20000$ data points from $\mu$ and train the score function s with denoising score matching (see Appendix D for diffusion models and Appendix H.1.1 for implementation details). In Figure 2 (left) we compare the evolution of the objective value of our approximation of (8) to the exact landing flow (10) for different noise levels $\sigma>0$ and dimension $n$. We observe that we can obtain objective values with cost lower than the best possible point in the training set and that the accuracy improves as $\sigma\to 0$.
Refer to caption Refer to caption
Figure 2: Objective value vs. flow time $t$ for the orthogonal manifold for $n=10$ (left, different $\sigma>0$) and for $n=20$ (right, $\sigma=0.05$). Here “exact landing” refers to equation 8 with exact operations $\texttt{v}=\pi$ and $\texttt{v}^{\prime}=\pi^{\prime}$.
@@ 6.2 Reference tracking via data-driven control
Under smoothness assumptions on the dynamics $\bm{f}$ and $\bm{g}$, the set $\mathcal{M}_{\operatorname{IO}}$, being a graph of a smooth map, is a (non-compact) embedded smooth submanifold of $(\mathbb{R}^{n_{u}})^{N_{h}}\times(\mathbb{R}^{n_{y}})^{N_{h}+1}$. The problem (14) is ubiquitous in receding horizon control applications such as model predictive control (MPC) and used for e.g. autonomous driving Vu et al. (2021), motion planning Cohen et al. (2020), optimizing HVAC system energy efficiency Serale et al. (2018) and inventory control Kostić (2009). In many of these applications the dynamics (13) governing the system are not known explicitly. Instead, in data-driven control Dörfler (2023a; b); Markovsky et al. (2023) one assumes that (13) is given implicitly by a finite number of measured input-output trajectories
@@
where the input $\bm{u}$ is persistently exciting Willems et al. (2005), e.g. given by (white) noise Ljung (1999). In particular the so-called system behavior manifold $\mathcal{M}_{\operatorname{IO}}$ Willems and Polderman (1997) is given by samples from a distribution $\mu$ on its in- and outputs and fits precisely into our framework of data-driven Riemannian optimization (1). A similar setup with observable state $\bm{y}=\bm{x}$ has been considered in the domain of reinforcement learning Janner et al. (2022). We test our proposed denoising Riemannian gradient descent on two classical systems from the control domain: The discretized double pendulum system and the unicycle car model LaValle (2006) (see Appendix H.2.1 detailed information on the systems and the particular choice of $\bm{r}$, $R$ and $Q$ in (15)), each on a horizon of $N_{h}=100$. To apply our proposed methods, we train a diffusion model (see Appendix H.2.2 for implementation details) on the measured trajectories $\mathcal{D}_{\operatorname{train}}$ and solve (14) via the denoising Riemannian gradient descent to obtain a solution $(\bm{u}^{*},\bm{y}^{*})$. As initial values we take the trajectories from the training set that minimize the objective cost $f$, i.e. $(\bm{u}_{0},\bm{y}_{0})=\operatorname*{arg\,min}_{(\bm{u},\bm{y})\in\mathcal{D}_{\operatorname{train}}}f(\bm{u},\bm{y})$. Note that, as seen in Section 5, in general the final iterate will not exactly lie on the input-output manifold, i.e. $(\bm{u}^{*},\bm{y}^{*})\notin\mathcal{M}_{\operatorname{IO}}$. To account for this deviation we back-test our generated input trajectory $\bm{u}^{*}$ by implementing it on the true underlying system (13) to obtain the real output $\bm{y}^{\operatorname{true}}$.
Results and discussion: In Figure 4 (in Appendix H.2.3) we depict the evolution of the objective value w.r.t. iteration count and in Figure 3 we depict the final optimizing trajectories $\bm{y}^{*}$ and $\bm{y}^{\operatorname{true}}$. We can observe that the error $\lVert\bm{y}^{*}-\bm{y}^{\operatorname{true}}\rVert$ is small, which shows that $(\bm{u}^{*},\bm{y}^{*})$ is close to the true system behavior $\mathcal{M}_{\operatorname{IO}}$. Moreover, we can see that the trajectory $\bm{y}^{\operatorname{true}}$ tracks the corresponding reference $\bm{r}$ much better than the train set minimum $\bm{y}_{0}$, which shows a generalization capability of our diffusion model. We from Figure 4 (left) that the current objective can depart from the true objective significantly. This is due to the iterates deviating from $\mathcal{M}_{\operatorname{IO}}$. In this example the algorithm (DRGD) recovers and we have found it to be robust w.r.t. moderate deviations from the manifold. Note that we have set a iteration budget of $N_{\operatorname{iter}}=3000$ and $N_{\operatorname{iter}}=2500$, respectively, while the objective is still decreasing. Accelerating the convergence of DRGD is left for future work.